# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ 5.1 Separation power: synthetic counterexamples
We construct a geometric expressiveness dataset based on counterexamples proposed by Li et al. [37] to evaluate the separation power of practical models. The synthetic dataset comprises 10 isolated and 7 combinatorial counterexamples. Each counterexample consists of a pair of highly symmetric point clouds (all of which are $\mathcal{D}$-symmetric described in Section 3) that are non-isomorphic but indistinguishable by DisGNN.
Table 2: The separation results on the constructed dataset.
@@
SchNet Vanilla Dis. DimeNet PaiNN MACE GeoNGNN
Isolated (10 cases) 0% 0% 100% 100% 100% 100%
Combin. (7 cases) 0% 0% 100% 100% 100% 100%
Results shown in Table 2 validate our theoretical findings. SchNet and DisGNN both fail to distinguish any counterexample, as expected. Conversely, the remaining models all effectively distinguish all pairs, which suggests that models integrating higher-order geometry or advanced aggregation strategies all possess the discrimination ability for existing carefully designed symmetric structures [37] despite potential numerical precision problems. We notice that we cannot conclude that PaiNN and MACE are theoretically complete from these results, and whether they are, with a finite degree of tensor order, still remains unclear and needs future investigation.
5.2 Molecule structure learning: revised MD17
We further evaluate GeoNGNN on the rMD17 dataset [8], which poses a substantial challenge to models’ geometric learning ability. This dataset encompasses trajectories from molecular dynamics simulations of several small molecules, and the objective is to predict the energy and atomic forces of a given molecule conformation, which contains all atoms’ positions and atomic numbers. The comprehensive results and comparison is shown in Table 1.
There are several important observations: 1) GeoNGNN exhibits substantial improvements over DisGNN. The ratio of the loss for GeoNGNN to that of DisGNN averages 1:8.6 and reaches a maximum of 1:34.2, validating the significance of capturing subgraph representations. 2) Though equally being E(3)-complete, DimeNet and GemNet show worse performance than GeoNGNN. This finding suggests that different model designs can exhibit quite different inductive biases, while learning more refined subgraph representations may be superior to directional message passing in certain molecule-relevant tasks. Although GeoNGNN demonstrates comparatively lower performance than another complete model 2F-DisGNN [37], in the subsequent subsection, we will show that on large point clouds (MD22 dataset) where 2F-DisGNN can hardly be applied due to computational infeasibility, GeoNGNN can still exhibit good performance while ensuring faster speed than existing SOTA models. 3) GeoNGNN shows competitive and even better performance in comparison to other well-designed state-of-the-art models. However, it does not consistently achieve SOTA performance, which may be due to its simple design, leveraging distance only, and the weak base GNN it adopts.
5.3 Supplementary experiments
In Appendix F, we show further evaluations of GeoNGNN, including: 1) Evaluation on MD22 [9], a dataset containing much larger molecules than MD17. Results show competitive performance and faster speed of GeoNGNN w.r.t. SOTA models such as MACE [3] and Equiformer [38]. 2) Evaluation on 3BPA dataset [33], where test sets comprise out-of-domain data, thereby emphasizing the significance of generalization ability. GeoNGNN shows competitive performance compared to Allegro [43], a well-designed equivariant geometric model. 3) Effect of subgraph size $r_{\text{sub}}$. Results imply that there exists a conflict between theoretical expressiveness and experimental performance, and striking a balance between them is important.
In Appendix H, we also defined, proved and evaluated GeoNGNN-C, an SE(3)-complete variant of GeoNGNN, designed by simply replacing the undirected distances in GeoNGNN with directed ones. GeoNGNN-C shows competitive or even better performance than SOTA models, such as ChiENN [19], on chirality-relevant datasets, even though it is not specifically designed for those tasks where additional symmetry, such as invariance under conformer-level transformation, could provide better inductive bias.
@@ 6 Conclusion and limitation
Conclusion
In this study, we thoroughly analyze a wide range of invariant models’ theoretical expressiveness. Specifically, we rigorously characterize the expressiveness of DisGNN, showcasing that all its unidentifiable cases exhibit $\mathcal{D}$-symmetry and have a measure of 0. We then establish a large family of E(3)-complete models, which encompasses GeoNGNN, geometric subgraph GNNs, as well as three established models - DimeNet, GemNet, and SphereNet. This contributes significantly to a comprehensive understanding of invariant geometric models. Moreover, the newly introduced geometric subgraph GNNs notably enlarge the design space of expressive geometric models. Experiments demonstrate the competitive performance of GeoNGNN, the simplest complete model, in comparison to geometric models utilizing high-order invariant and equivariant representations. Our experimental findings also suggest that models of equal completeness could exhibit distinct inductive biases on downstream tasks, thereby motivating further exploration to shift more focus from expressiveness towards generalization analysis.
Limitation
The E(3)-completeness for invariant models is under the conditions of fully connected graphs. However, the extent to which these models can exhibit high expressiveness on general sparse graphs remains an open question that needs further investigation. Additionally, future research is needed to characterize the expressiveness of vector models that rely solely on 1-order equivariant representations and adopt atom-level representations, such as PaiNN [52], which show promising experimental results but still lack theoretical expressiveness characterization.
@@ Appendix A Extended discussion of related work
To leverage the rich 3D geometric information contained in point clouds in permutation- and E(3)-invariant manner, early work Schütt et al. [53] integrated 3D Euclidean distance into the MPNN framework [24] and aggregate geometric information iteratively. Nevertheless, this model exhibits a restricted capacity to distinguish non-isomorphic point clouds even on fully-connected graphs as substantiated by Li et al. [37], Pozdnyakov & Ceriotti [46], Pozdnyakov et al. [47], Hordan et al. [26]. Subsequent invariant models [20, 21, 22, 23, 39, 56, 37] have endeavored to enhance their geometric expressiveness through the incorporation of carefully designed high-order geometric features, or adopting local spherical representations. These designs have greatly improved the models’ performance on downstream tasks, however, most of them lack theoretical guarantees of geometric completeness over the whole point cloud spaces.
Expressiveness of DisGNN
Prior work [47, 46, 37] first analyzed expressiveness of DisGNN through pair-wise separation of point clouds. However, they rely on individual carefully constructed counterexamples, which is not sufficient for rigorously characterize its expressiveness. Hordan et al. [26] establishes the completeness of DisGNN over a restricted subspace, which contains only point clouds with entirely distinct nodes, and therefore could be a coarse characterization.
General Geometric Expressiveness
Prior work Widdowson & Kurlin [58, 59], Kurlin [35] proposed complete invariants for point clouds, and continuous metrics for their comparison. However, these works involve no learnable parameters and do not directly correspond to neural function forms, restricting their applications in various downstream tasks. Joshi et al. [29] formally defines the expressiveness of geometric models through the separation of geometric graphs, extending similar concepts from traditional graph learning [60, 41, 42] to geometric settings. However, it does not fully characterize existing models’ expressiveness, such as revealing under what conditions can these models achieve E(3)-completeness. Moreover, it focuses on the expressiveness over geometric graphs, where node-level scalar and vector features are present, which do not exactly match our settings. Recently,  [37, 10, 26, 27] have proposed 2FWL-like invariant models with theoretical completeness, but such design space is restricted and whether weaker aggregation schemes, such as the scheme DimeNet [20] adopts, can achieve completeness remains unclear.
There are also lots of geometric completeness results in the literature, which, however, are somehow restricted. For example, Wang et al. [56] leverages the nearest neighboring node as a reference for computing higher-order geometric information. However, this leads to ambiguity when the point clouds are symmetric enough such that nodes have multiple nearest neighbors. This finally results in the completeness over a subspace of point clouds with only unsymmetric ones. Moreover, many frame-based methods project the coordinates or vector features onto local or global equivariant frames to ensure models’ Euclidean symmetry [57, 12, 48, 11]. Then, arbitrary powerful architectures, such as MLPs, are applied to these projected features without losing the Euclidean symmetry. Similarly, for most symmetric point clouds, which are exactly the corner cases that we aim to capture, the frames degenerate (can not be finitely determined). This implies that these methods may not be theoretically complete over all point clouds. We notice that there are several recent works that propose architectures based on weighted continuous frames [45, 31, 15] which could mitigate the degeneration and discontinuous problems. However, whether or under what conditions such models can achieve E(3)-completeness while maintaining exact symmetry over all point clouds remains unclear.
Finally, a significant proportion of geometric models leverage high-order SO(3) irreducible representations [55, 18, 3, 4, 6, 43], which come with the theoretical guarantee of universality [14]. However, the results [14] show that models need arbitrarily high-order spherical tensors for approximating high-order polynomials, and whether a finite order is sufficient still remains unclear.
Appendix B Extended analysis of DisGNN
B.1 The proper subset relation in Figure 1(b)

### v3

@@ 6.2 Separation Power on Synthetic Point Cloud Pairs
To construct hard-to-distinguish counterexamples, we develop a geometric expressiveness dataset based on the counterexamples proposed by Li et al. (2024). The synthetic dataset consists of 10 isolated and 7 combinatorial counterexamples. Each counterexample is composed of a pair of highly symmetric point clouds (all of which are $\mathcal{D}$-symmetric, as described in Section 4), which are non-isomorphic yet indistinguishable by DisGNN. We provide some examples in Figure 7, and please refer to Appendix G for further settings.
Table 1: Separation results on the constructed geometric expressiveness dataset. Models for which we have theoretically established completeness are highlighted in gray.
Invaraint Equivaraint
@@
SchNet DisGNN DimeNet SphereNet GemNet GeoNGNN PaiNN MACE
Isolated (10 cases) 0% 0% 100% 100% 100% 100% 100% 100%
Combined (7 cases) 0% 0% 100% 100% 100% 100% 100% 100%
Results are presented in Table 1. As shown, our established complete models can all distinguish these challenging pairs effectively, whereas DisGNN and SchNet cannot. This supports our theory and indicates that numerical precision is not impacting these complete models’ separation ability here. Interestingly, two equivariant models, PaiNN (Schütt et al., 2021) and MACE (Batatia et al., 2022) that leverages vectors and high-order tensors respectively, also effectively distinguish all the pairs. This prompts further question of whether they are also complete, especially under sparse connections or finite tensor orders, and requires further investigation.
@@ 7 Conclusion and Limitation
Conclusion. In this study, we thoroughly analyze a wide range of invariant models’ theoretical expressiveness under fully-connected condition. Specifically, we rigorously characterize the expressiveness of DisGNN, showcasing that all its unidentifiable cases exhibit $\mathcal{D}$-symmetry and have a measure of 0. We then establish a large family of E(3)-complete models, which encompasses GeoNGNN, geometric subgraph GNNs, as well as three established models - DimeNet, GemNet, and SphereNet. This contributes significantly to a comprehensive understanding of invariant geometric models. Moreover, the newly introduced geometric subgraph GNNs notably enlarge the design space of expressive geometric models. Experiments further validate our theoretical findings.
Limitation. The rigorous E(3)-completeness characterization for invariant models is under the conditions of fully connected graphs, which can be limited and have been throughout discussed in Section 5.4. The extent to which invariant models can exhibit high expressiveness on general sparse graphs remains an open question that needs further investigation. Additionally, future research is needed to characterize the expressiveness of vector models that rely solely on 1-order equivariant representations and adopt atom-level representations, such as PaiNN (Schütt et al., 2021), which show promising experimental results but still lack theoretical expressiveness characterization.
Acknowledgment
This work is supported by the National Natural Science Foundation of China (62276003).
@@ Appendix A Extended discussion of related work
To leverage the rich 3D geometric information contained in point clouds in permutation- and E(3)-invariant manner, early work Schütt et al. (2018) integrated 3D Euclidean distance into the MPNN framework (Gilmer et al., 2017) and aggregate geometric information iteratively. Nevertheless, this model exhibits a restricted capacity to distinguish non-isomorphic point clouds even on fully-connected graphs as substantiated by Li et al. (2024); Pozdnyakov & Ceriotti (2022); Pozdnyakov et al. (2020); Hordan et al. (2024b). Subsequent invariant models (Gasteiger et al., 2019; 2020; 2021; 2022; Liu et al., 2021; Wang et al., 2022; Li et al., 2024) have endeavored to enhance their geometric expressiveness through the incorporation of carefully designed high-order geometric features (Gasteiger et al., 2019; 2020; 2021; 2022), adopting local spherical representations (Wang et al., 2022) or higher-order distance features (Li et al., 2024). These designs have greatly improved the models’ performance on downstream tasks, however, most of them lack theoretical guarantees of geometric completeness over the whole point cloud spaces.
Appendix B Proof of Main Conclusions in Main Body
B.1 Measure of Symmetric and Unidentifiable Point Clouds
In this subsection, we aim to prove that the Lebesgue measure on $\mathbb{R}^{n\times 3}$ of the $\mathcal{C}$-symmetric point cloud set, the $\mathcal{D}$-symmetric point cloud set, and the unidentifiable point cloud set are all zero.
Theorem 4.3 (Unidentifiable set of DisGNN has measure zero) The Lebesgue measure on $\mathbb{R}^{n\times 3}$ of the $\mathcal{C}$-symmetric, $\mathcal{D}$-symmetric, and unidentifiable point cloud sets is zero.
We first show that the $\mathcal{C}$-symmetric point cloud set has measure zero. We denote the set containing all $\mathcal{C}$-symmetric point clouds as $\mathbb{R}^{n\times 3}_{\mathcal{C}}$. Note that
$\mathbb{R}^{n\times 3}_{\mathcal{C}}\subseteq\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}:=\{P\in\mathbb{R}^{n\times 3}\mid\exists i\neq j\in[n]\text{ such that }\text{cond}_{i,j}\text{ holds}\},$
where
$\text{cond}_{i,j}:=\|p_{i}-\text{mean}_{k}(p_{k})\|^{2}-\|p_{j}-\text{mean}_{k}(p_{k})\|^{2}=0,$
which is essentially a non-trivial polynomial equality constraint. Here, $\text{mean}_{k}(p_{k})$ calculates the geometric center of the point cloud, and $\|p_{i}-\text{mean}_{k}(p_{k})\|$ is essentially the distance between $i$ and the geometric center. The set $\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}$ is a superset of $\mathbb{R}^{n\times 3}_{\mathcal{C}}$, which can be validated as follows:
Assume some element $P$ belongs to $\mathbb{R}^{n\times 3}_{\mathcal{C}}$ but not to $\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}$. Then, we have
$\forall i\neq j\in[n],\quad\text{cond}_{i,j}\text{ for $P$ does not hold}.$
This essentially means that all the nodes are embedded with different features after applying the $\mathcal{C}$ algorithm. Then, it is obvious that $P$ is $\mathcal{C}$-asymmetric, which contradicts the assumption that $P\in\mathbb{R}^{n\times 3}_{\mathcal{C}}$. Hence, we conclude that
$\mathbb{R}^{n\times 3}_{\mathcal{C}}\subseteq\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}.$
The set $\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}$ defines an algebraic manifold with a non-trivial polynomial equality constraint, thus having dimension at most $13n-$ 1. Therefore, according to (Mityagin, 2015; Hordan et al., 2024b), the set $\mathbb{R}^{n\times 3}_{\text{super }\mathcal{C}}$ has measure zero, which implies that $\mathbb{R}^{n\times 3}_{\mathcal{C}}$ also has measure zero.
According to Proposition C.1, since $\mathbb{R}^{n\times 3}_{\mathcal{C}}$ is the largest set under consideration, it follows that both the $\mathcal{D}$-symmetric point cloud set and the unidentifiable point cloud set also have measure zero.
B.2 Preparation for Reconstruction Proof
Definition B.1.
For example, consider the space $\mathbb{P}=\mathbb{R}^{n\times 3}$ comprising all point clouds of size $n$. Let $f$ denote an encoder that computes $f(P)=(\sum_{i=1}^{n}p_{i},n)$, which represents the sum of the nodes’ coordinates and the number of nodes, while $g$ being an encoder calculating the geometric center of $P$, it follows that $f(P)\to g(P)$, since we can calculate the geometric center from the sum of the nodes’ coordinates and the number of nodes. In the following analysis, we consider $\mathbb{P}=\mathbb{R}^{n\times 3}$ by default. Notice that derivation exhibits transitivity: if $f_{1}\to f_{2}$ and $f_{2}\to f_{3}$, it follows that $f_{1}\to f_{3}$.
Intuitively, “derive” is a concept describing the relation of two properties, i.e., whether property $f(P)$ contains all the information needed to calculate $g(P)$. Obviously, if an encoder $f$ can embed all the information needed for reconstructing a given point cloud, then it is E(3)-complete: